OpenStudy (anonymous):

what is the cost in dollars for the nails used to build a fence 125 meters long if it requires 30 nails per meter? Assume that 40 nails are sold per box at a cost of 75 cents per box

OpenStudy (anonymous):

oh and it has to be solved the factor label or dimensional analysis way

OpenStudy (anonymous):

Can explain what factor label and dimensional analysis. I just solved it n got $7.50

OpenStudy (anonymous):

oh its like converting units you can google it

OpenStudy (anonymous):

30nails Per Meter \[\frac{ 30nails }{ 1meter }=\frac{ ? }{125meters }\] Multiply both sides by 125Meters and meters cancel out \[125*\frac{ 30nails }{ 1 }=\frac{ ? }{125}*125\] We get 3750 nails are needed To find out how many boxes we have to divide by 40 \[\frac{ 3750 nails }{ ? }= \frac{ 40 nails }{ 1box }=\frac{ 3750 nails }{ 40 nails }*1box=93.5boxes\] must rounded to nearest whole number \[\approx 94 boxes\] \[94*$.75=$70.50\] So correction on my first attempt

OpenStudy (anonymous):

thanks thats exactly what i was trying to get too

OpenStudy (anonymous):

you're welcome i put 4 nails on my first attempt instead of 40

OpenStudy (anonymous):

oh thats probably why the first one looked weird